OpenStudy (anonymous):

How can I solve sin 70 cos 50 + sin 50 cos 70 without using a calculator?

OpenStudy (anonymous):

this is the formula for the sine of sum two angles: sin(A+B) = sinAcosB + sinBcosA notice that 120 = 70 + 50. so sin70cos50 + sin50cos70 = sin(70+50) = sin(120) = sqrt3/2

OpenStudy (anonymous):

Oh, I get it. Thanks a lot